OpenStudy (anonymous):

FAN AND MEDAL, HELP PLEASE? =) multiplication involving binomials and trinomials in one variable; (7u+3)(2u^2-6u-1) simplify.

OpenStudy (anonymous):

7u+3(2u^2)= 14u^3+6u^2 7u+3(-6u)= -42u^2-18u 7u+3(-1)= -7u-3 14u^3+6u^2-42u^2-18u-7u-3 14u^3-36u^2-25u-3 ?

OpenStudy (anonymous):

You have to do long division to complete these. Ex. (2x + 1) (3x^2 + 2x - 1) 2x + 1 | (4X^2 + 2x - 1) 1.) 2x times what gives you 4x^2? = 2x So put 4x^2 under the 4x^2. (they'll cancel out when subtracting). 2.) Now, what is 2x times 1? = 2x. So put + 2x beside the number you just wrote. 3.) Subtract them. The first numbers should cancel out, then you'll have a remainder with the other one. Then drop down the 2x from the trinomial: Just repeat that until there aren't anymore numbers to drop down.

OpenStudy (anonymous):

thank you! i'll try it haha I've never been good at math =/